OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (anonymous):

you could check each ordered pair and see which one works

OpenStudy (anonymous):

or start with \[x + 3y = -1 \\2x + 2y = 6 \] multiply the first equation by -2 and get \[-2x-6y=2\\2x+2y=6\] then add them up to get \[5y=5\] telling you \(y=1\)

OpenStudy (anonymous):

whoa that was wrong!

OpenStudy (anonymous):

add them up and get \[-4y=8\] and so \[y=-2\]

OpenStudy (anonymous):

Thank you(: So the answer would be D?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you help me with one more ?